# Rethinking Multimodal Learning from the Perspective of Mitigating Classification Ability Disproportion

### v1

@@ Abstract
Although multimodal learning (MML) has garnered remarkable progress, the existence of modality imbalance hinders multimodal learning from achieving its expected superiority over unimodal models in practice. To overcome this issue, mainstream multimodal learning methods have placed greater emphasis on balancing the learning process. However, these approaches do not explicitly enhance the classification ability of weaker modalities, leading to limited performance promotion. By designing a sustained boosting algorithm, we propose a novel multimodal learning approach to dynamically balance the classification ability of weak and strong modalities. Concretely, we first propose a sustained boosting algorithm in multimodal learning by simultaneously optimizing the classification and residual errors using a designed configurable classifier module. Then, we propose an adaptive classifier assignment strategy to dynamically facilitate the classification performance of weak modality. To this end, the classification ability of strong and weak modalities is expected to be balanced, thereby mitigating the imbalance issue. Empirical experiments on widely used datasets reveal the superiority of our method through comparison with various state-of-the-art (SoTA) multimodal learning baselines.
@@ 1 Introduction
In recent years, multimodal learning [30, 31, 39, 45, 19] has received growing attention for its ability to effectively integrate heterogeneous information. As extra information from multimodal data can be utilized, MML is expected to achieve better performance compared with unimodal approaches. However, contrary to expectations, MML has been surprisingly shown to underperform compared to unimodal ones in certain scenarios [37, 31].
Figure 1: Comparison with naive MML, gradient boosting based MML (MML w/ GB), G-Blend [37], and Ours on CREMAD dataset. We find that enhancing the classification performance of the weak modality narrows the performance gap between the two modalities and improves overall performance.
@@
The root of this problem lies in the existence of the modality imbalance [37]. Concretely, different modalities in a joint-training paradigm typically converge at different speeds [31, 40]. The faster-converging modality, i.e., strong modality [41], tends to achieve higher performance, while the weak modality performs poorly. Subsequently, this disproportion in classification ability often leads to modality imbalance [37], ultimately resulting in lower performance.
@@
Researchers have explored the modality imbalance issue from various perspectives in multimodal learning [37, 31, 45]. Given the inconsistent learning progress between strong and weak modalities, a natural idea [37, 31, 26, 43] is to manually intervene in their learning processes to achieve rebalancing. Another type of method is to bridge the information gap between modality training phases and enhance the interaction between different modalities during training. To be specific, impressive works [45, 12] such as MLA [45], ReconBoost [19] and DI-MML [12] focus on bridging the learning gap of different modalities through injecting the optimization information between modalities.
@@
Although the above methods can rebalance multimodal learning, they focus more on balancing the learning process while failing to enhance the classification ability explicitly. Compared to weaker modalities, stronger modalities typically yield more robust classifiers due to their more sufficient information [41]. Is there a way to directly improve the performance of weak classifiers to balance the classification performance between strong and weak modalities? A natural choice is boosting [13, 14], which utilizes the ensemble technique to enhance the ability of the weak classifier. We conduct a toy experiment to illustrate this idea on CREMAD dataset [4], where the classifier of weak modality is enhanced by the gradient boosting [14]. The results in Figure 1 present the comparison among naive MML, a model learning adjustment-based MML approach (G-Blend [37]), and gradient boosting-based MML (MML w/ GB). For MML w/ GB, we apply the gradient boosting algorithm to further improve the trained video model using naive MML, while keeping the audio model fixed. We can find that the classification gap between video and audio modalities of naive MML and G-Blend is relatively large. More importantly, for MML w/ GB, the accuracy of audio modality remains unchanged, but the accuracy of video is greatly improved, leading to the improvement of overall accuracy.
@@
According to the aforementioned observations, in this paper, we propose a novel multimodal learning approach by designing a sustained boosting algorithm to facilitate the classification ability of the weaker modality. Concretely, we first design a configurable classification module, called the configurable classifier. This module takes features extracted by the encoder as input and provides predictions for the given data. We propose a sustained boosting algorithm by using this module as a basic classifier. Then, we utilize OGM [31] score to monitor the learning status during joint training, and further propose an adaptive classifier assignment (ACA) strategy to adjust the classifier of weak modality. To this end, we can enhance the classification ability for the weak modality, thereby rebalancing the classification ability of strong and weak modalities. In Figure 1, we present the classification enhancement results of our method (Ours). We can find that the performance of our method outperforms that of MML w/ GB thanks to the sustained boosting and adaptive classifier assignment strategy. Furthermore, it is worth mentioning that ReconBoost [19] also employs the gradient boosting algorithm for multimodal learning. However, unlike our approach, ReconBoost uses gradient boosting to iteratively learn complementary information across modalities. Our main contributions are outlined as follows:
  • We propose a sustained boosting algorithm in MML to simultaneously minimize the classification and residual errors based on a designed configurable classifier module.
@@
  • Based on the learning status, we propose a novel adaptive classifier assignment strategy to dynamically enhance the classification ability of weak modality, thus rebalancing the classification ability of all modalities.
@@
  • We conduct comprehensive experiments to verify the effectiveness of our approach. Results demonstrate that our approach can outperform SoTA baselines to achieve the best performance by a large margin.
@@ 2 Related Work
2.1 MML under Imbalanced Scenario
The goal of multimodal learning [42, 22, 30, 34, 20] is to fuse the multimodal information from diverse sensors. Compared to unimodal methods, MML can mine data information from different perspectives, thus the performance of multimodal learning should be better [28, 33, 18, 15]. However, due to heterogeneity of multimodal data, multimodal learning often encounters imbalance problems [37, 21] in practice, leading to performance degeneration of MML.
Early pioneering works [37, 31, 11, 16] focus more on adaptively adjusting the learning procedure for different modalities. Representative approaches in this category employ different learning strategies, e.g., gradient modulation [31, 26] and learning rate adjustment [43], to rebalance the learning of weak and strong modalities. Other approaches including MLA [45], DI-MML [12], ReconBoost [19] and MAIE [23] take a different path, focusing on enhancing the interaction between modalities to address the modality imbalance problem. For example, MLA [45] designs an alternating algorithm to train different modalities iteratively. During the training phase, the interaction is enhanced by transferring the learning information between different modalities. ReconBoost [19] balances modality learning by leveraging gradient boosting to capture information from other modalities during interactive learning.
@@ 2.2 Boosting Method
Boosting algorithm [24, 13, 14, 8, 27, 32, 9] is one of the most important algorithms in ensemble learning. The core idea of boosting is to integrate multiple learners to create a strong learner. Adaboost [13], one of the earliest boosting algorithms, adjusts the weights of incorrectly classified data points, giving more attention to the harder-to-classify examples in each iteration. Gradient boosting [14], on the other hand, builds models in a stage-wise fashion, minimizing a loss function through gradient descent.
@@
Figure 2: The framework of our proposed method. We utilize the video and audio modalities as examples, with the numbers of video and audio classifiers denoted by $n^{v}$ and $n^{a}$, respectively.
3 Problem Definition
3.1 Notation
In this paper, we use boldface lowercase letters like ${\bf z}$ to denote vectors. The symbol $\odot$ is used to denote the Hadamard product. We use $\|\cdot\|_{2}$ to denote the $L_{2}$ norm of vectors. Furthermore, $\delta(\cdot)$ denotes the indicator function, i.e., $\delta(true)=1$, otherwise $\delta(false)=0$. $\mathtt{min}(\cdot)$ denotes the function that returns the minimum value. $\mathtt{mod}(a,b)$ returns the remainder after division of $a$ by $b$.
@@ 3.2 Multimodal Learning
Assume that we have $N$ data points, each of which has audio and video modalities. Without loss of generality, we use ${\bf X}=\{({\bf x}_{i}^{a},{\bf x}^{v}_{i})\}_{i=1}^{N}$ to denote the multimodal data, where ${\bf x}_{i}^{a}$ and ${\bf x}_{i}^{v}$ denote the $i$-th data point of audio and video, respectively. In addition, we are also given a category labels set ${\bf Y}=\{\;{\bf y}_{i}~{}|~{}{\bf y}_{i}\in\{0,1\}^{K}\}_{i=1}^{N}$, where $K$ denotes the number of category labels. Given the above training information ${\bf X}$ and ${\bf Y}$, the goal of multimodal learning is to train a model to fuse the multimodal information and predict its category label as accurately as possible.
4 Methodology
In this section, we present our method in detail. The architecture of our method is shown in Figure 2, where the audio and video modalities are used as an example for illustration.
@@ 4.1 Sustained Boosting
For the sake of simplicity, we use superscript $r$ to indicate the module corresponding to a specific modality in this section, where $r\in\{a,v\}$. With the rapid growth of deep learning, representative MML approaches [30, 37, 11, 26] have adopted deep neural network (DNN) for multimodal learning. Following these methods, we also utilize DNN to construct our models. Specifically, we use $\phi^{r}(\cdot)$ to denote encoders. Then the features can be calculated by ${\bf u}^{r}=\phi^{r}({\bf x}^{r};\theta^{r})$, where $\theta^{r}$ denotes the encoder parameters. Then, the prediction of given data can be calculated by a classifier $\psi^{r}(\cdot)$: ${\bf p}^{r}=\psi^{r}({\bf u}^{r};\Theta^{r}),$ where $\Theta^{r}$ denotes the parameters of the classifier. Based on ${\bf p}^{r}$ and its ground-truth, the objective function can be written as:
${\mathcal{L}}_{\textit{CE}}({\bf X}^{r},{\bf Y})=\frac{1}{N}\sum_{i=1}^{N}\ell({\bf p}^{r}_{i},{\bf y}_{i})=-\frac{1}{N}\sum_{i=1}^{N}{\bf y}_{i}^{\top}\log({\bf p}^{r}_{i}),$ (1)
where $\Phi^{r}\triangleq\{\theta^{r},\Theta^{r}\}$ denotes the parameters to be learned, ${\bf X}^{r}\triangleq\{{\bf x}^{r}_{i}\}_{i=1}^{N}$ and $\ell(\cdot)$ denotes the cross entropy loss.
@@
By training the model for each modality based on objective function (1), we can obtain multiple individual classifiers. Due to the existence of strong and weak modalities [41], these classifiers exhibit different classification abilities. Hence, we can employ boosting technique [14] to improve the classification ability of weak modality.
@@
Concretely, assuming the classification performance of the $r$-th modality requires improvement, we first apply the gradient boosting algorithm to train $n$ classifiers for the $r$-th modality. Since feature extraction focuses on common patterns, we set the encoders of all classifiers to be shared. Then the $j$-th classifier can be defined as: $\Phi^{r}_{t}\triangleq\{\theta^{r},\Theta^{r}_{t}\},t\in\{1,\cdots,n\}$. In practice, we adopt multiple fully-connected layers and nonlinear activation rectified linear unit (ReLU) [1] to construct our classification module. This module called the configurable classifier, is relatively independent and can be adjusted based on the classification ability. Furthermore, we adopt the shared head structure commonly used in MML [45, 23, 7] to strengthen the interaction between weak and strong modalities during training.
@@
Inspired by gradient boosting [14], the classification ability can be facilitated through minimizing the residual error introduced by previous classifiers. Concretely, when we learn $t$-th classifier, the residual labels are defined as:
$\hat{\bf y}^{r}_{it}={\bf y}_{i}-\lambda\sum_{j=1}^{t-1}{\bf y}_{i}\odot{\bf p}^{r}_{ij},$
@@
where $\lambda\in[0,1]$ is used to soften hard labels [35] and we utilize ${\bf y}_{i}$ to mask non ground-truth labels to ensure the non-negativity of residual labels. Then the objective function can be defined as follows:
$\epsilon({\bf x}^{r}_{i},{\bf y}_{i},t)=\ell\big{(}{\bf p}^{r}_{it},\hat{\bf y}^{r}_{it}\big{)},$ (2)
@@
where ${\bf p}^{r}_{it}$ denotes the prediction obtained by $t$-th classifier for $i$-th data point. Since we utilize a shared encoder, the encoder will be updated when training the $t$-th classifier. Therefore, other classifiers must be updated simultaneously to prevent performance degradation. The corresponding objective can be formed as:
$\epsilon_{o}({\bf x}^{r}_{i},{\bf y}_{i},t)$ $=\ell\left({\bf p}^{r}_{it}+\sum_{j=1}^{t-1}{\bf p}^{r}_{ij},{\bf y}_{i}\right)=\ell\left(\sum_{j=1}^{t}{\bf p}^{r}_{ij},{\bf y}_{i}\right).$ (3)
@@
Meanwhile, we have to ensure the first $t-1$ classifiers are well-trained. Hence, we define the following objective for $t-1$ classifiers:
$\epsilon_{p}({\bf x}^{r}_{i},{\bf y}_{i},t)$ $=\ell\left(\sum\nolimits_{j=1}^{t-1}{\bf p}^{r}_{ij},{\bf y}_{i}\right).$ (4)
@@
$L({\bf x}^{r}_{i},{\bf y}_{i},t)=\epsilon({\bf x}^{r}_{i},{\bf y}_{i},t)+\epsilon_{o}({\bf x}^{r}_{i},{\bf y}_{i},t)+\epsilon_{p}({\bf x}^{r}_{i},{\bf y}_{i},t).$ (5)
Unlike traditional gradient boosting [14], our method sustainedly minimizes classification and residual errors by optimizing (5). Then the overall loss of sustained boosting can be formed as:
${\mathcal{L}}_{\textit{SUB}}({\bf X}^{r},{\bf Y};\Phi^{r})=$ $\frac{1}{N}\sum\nolimits_{i=1}^{N}L({\bf x}^{r}_{i},{\bf y}_{i},n).$ (6)
Algorithm 1 ACA algorithm.
1:OGM score ${\bf s}_{t}$.
2:Classification weight ${\boldsymbol{\omega}}_{t}$ and assignment decision.
3:INIT initialize assignment status as false.

### v2

@@ Abstract
Multimodal learning (MML) is significantly constrained by modality imbalance, leading to suboptimal performance in practice. While existing approaches primarily focus on balancing the learning of different modalities to address this issue, they fundamentally overlook the inherent disproportion in model classification ability, which serves as the primary cause of this phenomenon. In this paper, we propose a novel multimodal learning approach to dynamically balance the classification ability of weak and strong modalities by incorporating the principle of boosting. Concretely, we first propose a sustained boosting algorithm in multimodal learning by simultaneously optimizing the classification and residual errors. Subsequently, we introduce an adaptive classifier assignment strategy to dynamically facilitate the classification performance of the weak modality. Furthermore, we theoretically analyze the convergence property of the cross-modal gap function, ensuring the effectiveness of the proposed boosting scheme. To this end, the classification ability of strong and weak modalities is expected to be balanced, thereby mitigating the imbalance issue. Empirical experiments on widely used datasets reveal the superiority of our method through comparison with various state-of-the-art (SOTA) multimodal learning baselines. The source code is available at https://github.com/njustkmg/NeurIPS25-AUG.
@@ 1 Introduction
In recent years, multimodal learning MMDL:conf/icml/NgiamKKNLN11 ; MML:conf/ijcai/YangWZX019 ; SMV:conf/cvpr/YakeRZD24 ; LFM:conf/nips/0074WJ024 ; ReconBoost:conf/icml/CongHua24 has received growing attention for its ability to effectively integrate heterogeneous information. As extra information from multimodal data can be utilized, multimodal learning is expected to achieve better performance compared with unimodal approaches. However, contrary to expectations, multimodal learning has been surprisingly shown to underperform compared to unimodal ones in certain scenarios OGR-GB:conf/cvpr/WangTF20 ; OGM:conf/cvpr/PengWD0H22 ; AMSS:journals/pami/YangPJXT25 .
@@
The root of this problem lies in the existence of the modality imbalance OGR-GB:conf/cvpr/WangTF20 . Concretely, different modalities in a joint-training paradigm typically converge at different speeds OGM:conf/cvpr/PengWD0H22 ; nvGesture:conf/icml/WuJCG22 . The faster-converging modality, i.e., strong modality ARM:conf/ijcai/YangYZJ15 , tends to achieve higher performance, while the weak modality performs poorly. Subsequently, this disproportion in classification ability often leads to modality imbalance OGR-GB:conf/cvpr/WangTF20 , ultimately resulting in lower performance.
@@
Researchers have explored the modality imbalance issue from various perspectives in multimodal learning OGR-GB:conf/cvpr/WangTF20 ; OGM:conf/cvpr/PengWD0H22 ; MLA:conf/cvpr/ZhangYBY24 . Given the inconsistent learning progress between strong and weak modalities, a natural idea OGR-GB:conf/cvpr/WangTF20 ; OGM:conf/cvpr/PengWD0H22 ; AGM:conf/iccv/LiLHLLZ23 ; MSLR:conf/acl/YaoM22 is to manually intervene in their learning processes to achieve rebalancing. Another type of method is to bridge the information gap between modality training phases and enhance the interaction between different modalities during training. To be specific, impressive works MLA:conf/cvpr/ZhangYBY24 ; DI-MML:conf/mm/FanXWLG24 such as MLA MLA:conf/cvpr/ZhangYBY24 , ReconBoost ReconBoost:conf/icml/CongHua24 and DI-MML DI-MML:conf/mm/FanXWLG24 focus on bridging the learning gap of different modalities through injecting the optimization information between modalities.
@@
Although the above methods can rebalance multimodal learning, they focus more on balancing the learning process while failing to enhance the classification ability explicitly. Compared to weaker modalities, stronger modalities typically yield more robust classifiers due to their more sufficient information ARM:conf/ijcai/YangYZJ15 . Is there a way to directly improve the performance of weak classifiers to balance the classification performance between strong and weak modalities? A natural choice is boosting Adaboost:conf/eurocolt/FreundS95 ; GradientBoosting:journal/AS/FriedmanJ , which utilizes the ensemble technique to enhance the ability of the weak classifier. We conduct a toy experiment to illustrate this idea on CREMAD dataset CREMAD:journals/taffco/CaoCKGNV14 , where the classifier of weak modality is enhanced by the gradient boosting GradientBoosting:journal/AS/FriedmanJ . The results in Figure 1 present the comparison among naive MML, a model learning adjustment-based MML approach (G-Blend OGR-GB:conf/cvpr/WangTF20 ), and gradient boosting-based MML (MML w/ GB). For MML w/ GB, we apply the gradient boosting algorithm to further improve the trained video model using naive MML, while keeping the audio model fixed. We can find that the classification gap between video and audio modalities of naive MML and G-Blend is relatively large. More importantly, for MML w/ GB, the accuracy of audio modality remains unchanged, but the accuracy of video is greatly improved, leading to the improvement of overall accuracy. This demonstrates the feasibility and effectiveness of using boosting to balance the classification ability of strong and weak modalities in mitigating modality imbalance.
Figure 1: Comparison with naive MML, gradient boosting based MML (MML w/ GB), G-Blend OGR-GB:conf/cvpr/WangTF20 , and Ours on CREMAD dataset. We find that enhancing the classification performance of the weak modality narrows the performance gap between the two modalities and improves overall performance.
@@
According to the aforementioned observations, in this paper, we propose a novel multimodal learning approach by designing a sustained boosting algorithm to facilitate the classification ability of the weaker modality. Concretely, we first propose a sustained boosting algorithm by jointly optimizing the classification loss and residual error, aiming to enhance the classification performance of the weak modality. The algorithm takes features extracted by the encoder as input and provides predictions for the given data. Then, we employ the confident score OGM:conf/cvpr/PengWD0H22 to monitor the learning status during joint training, and further propose an adaptive classifier assignment (ACA) strategy to adjust the classifier of weak modality. To this end, we can enhance the classification ability for the weak modality, thereby rebalancing the classification ability of strong and weak modalities. Meanwhile, we theoretically show that, under the boosting algorithm framework, the gap between different losses is provably convergent. In Figure 1, we present the classification enhancement results of our method (Ours). We can find that the performance of our method outperforms that of MML w/ GB thanks to the sustained boosting and adaptive classifier assignment strategy. Furthermore, it is worth mentioning that ReconBoost ReconBoost:conf/icml/CongHua24 also employs the gradient boosting algorithm for multimodal learning. However, unlike our approach, ReconBoost uses gradient boosting to iteratively learn complementary information across modalities. Our main contributions are outlined as follows:
  • A novel sustained boosting algorithm in MML is proposed. This algorithm aims to simultaneously minimize the classification and residual errors to facilitate the classification ability of the weak modality.
@@
  • A novel adaptive classifier assignment strategy is proposed to dynamically enhance the classification ability of weak modality based on the learning status, thus rebalancing the classification ability of all modalities.
  • We theoretically analyze the impact of the boosting algorithm on the loss gap between different modalities and prove its convergence.
@@
  • Experiments reveal that our approach can outperform SOTA baselines to achieve the best performance by a large margin on widely used datasets.
@@ 2 Related Work
2.1 Rebalanced Multimodal Learning
The goal of multimodal learning MML:conf/ijcai/YangWZX019 ; MMDL:conf/icml/JiaYXCPPLSLD21 ; MMDL:conf/icml/NgiamKKNLN11 ; MMDL:conf/cvpr/SinghHGCGRK22 ; MML:conf/nips/HuangDXCZH21 is to fuse the multimodal information from diverse sensors. Compared to unimodal methods, multimodal learning can mine data information from different perspectives, thus the performance of multimodal learning should be better MML:conf/cvpr/LvCHDL21 ; MML:conf/nips/SimonyanZ14 ; MML:conf/cvpr/HuLL16 ; MML:conf/cvpr/GaoOGT20 . However, due to heterogeneity of multimodal data, multimodal learning often encounters imbalance problems OGR-GB:conf/cvpr/WangTF20 ; ModalCompetition:conf/icml/HuangLZYH22 in practice, leading to performance degeneration of multimodal learning.
Early pioneering works OGR-GB:conf/cvpr/WangTF20 ; OGM:conf/cvpr/PengWD0H22 ; PMR:conf/cvpr/Fan0WW023 ; MML:conf/cvpr/Ge0G00AILZ23 focus more on adaptively adjusting the learning procedure for different modalities. Representative approaches in this category employ different learning strategies, e.g., gradient modulation OGM:conf/cvpr/PengWD0H22 ; AGM:conf/iccv/LiLHLLZ23 and learning rate adjustment MSLR:conf/acl/YaoM22 , to rebalance the learning of weak and strong modalities. Other approaches including MLA MLA:conf/cvpr/ZhangYBY24 , ReconBoost ReconBoost:conf/icml/CongHua24 and IGM IGM:conf/ijcai/JiangCY25 take a different path, focusing on enhancing the interaction between modalities to address the modality imbalance problem. For example, MLA MLA:conf/cvpr/ZhangYBY24 designs an alternating algorithm to train different modalities iteratively. During the training phase, the interaction is enhanced by transferring the learning information between different modalities. ReconBoost ReconBoost:conf/icml/CongHua24 balances modality learning by leveraging gradient boosting to capture information from other modalities during interactive learning. IGM IGM:conf/ijcai/JiangCY25 employs a flat gradient modification strategy to enhance the interactive multimodal learning.
@@ 2.2 Boosting Method
Boosting algorithm Adaboost:conf/eurocolt/FreundS95 ; GradientBoosting:journal/AS/FriedmanJ ; Boost:conf/cvpr/LiuHMT14 ; Boost:conf/icml/CortesMS14 ; Boost:conf/nips/ProkhorenkovaGV18 is one of the most important algorithms in ensemble learning. The core idea of boosting is to integrate multiple learners to create a strong learner. Adaboost Adaboost:conf/eurocolt/FreundS95 , one of the earliest boosting algorithms, adjusts the weights of incorrectly classified data points, giving more attention to the harder-to-classify examples in each iteration. Gradient boosting GradientBoosting:journal/AS/FriedmanJ , on the other hand, builds models in a stage-wise fashion, minimizing a loss function through gradient descent. It iteratively refines the overall model by fitting the negative gradient of the loss function GB:journals/finr/NatekinK13 with respect to the model’s predictions.
@@
3 Methodology
@@ 3.1 Multimodal Learning
Assume that we have $N$ data points, each of which has audio and video modalities. Without loss of generality, we use ${\boldsymbol{X}}=\{({\boldsymbol{x}}_{i}^{a},{\boldsymbol{x}}^{v}_{i})\}_{i=1}^{N}$ to denote the multimodal data, where ${\boldsymbol{x}}_{i}^{a}$ and ${\boldsymbol{x}}_{i}^{v}$ denote the $i$-th data point of audio and video, respectively. In addition, we are also given a category labels set ${\boldsymbol{Y}}=\{\;{\boldsymbol{y}}_{i}~|~{\boldsymbol{y}}_{i}\in\{0,1\}^{K}\}_{i=1}^{N}$, where $K$ denotes the number of category labels. Given the above training information $\boldsymbol{X}$} and $\boldsymbol{Y}$}, the goal of multimodal learning is to train a model to fuse the multimodal information and predict its category label as accurately as possible.
For the sake of simplicity, we use superscript $o$ to indicate the module corresponding to a specific modality in this section, where $o\in\{a,v\}$. With the rapid growth of deep learning, representative MML approaches MMDL:conf/icml/NgiamKKNLN11 ; OGR-GB:conf/cvpr/WangTF20 ; PMR:conf/cvpr/Fan0WW023 ; AGM:conf/iccv/LiLHLLZ23 have adopted deep neural network (DNN) for multimodal learning. Following these methods, we also utilize DNN to construct our models. Specifically, we use $\phi^{o}(\cdot)$ to denote encoders. Then the features can be calculated by ${\boldsymbol{u}}^{o}=\phi^{o}({\boldsymbol{x}}^{o};\theta^{o})$, where $\theta^{o}$ denotes the encoder parameters. Then, the prediction of given data can be calculated by a classifier $\psi^{o}(\cdot)$: ${\boldsymbol{p}}^{o}=\psi^{o}({\boldsymbol{u}}^{o};\Theta^{o}),$ where $\Theta^{o}$ denotes the parameters of the classifier. Based on ${\boldsymbol{p}}^{o}$ and its ground-truth, the objective function can be written as:
${\mathcal{L}}_{\mathrm{CE}}({\boldsymbol{X}}^{o},{\boldsymbol{Y}})=\frac{1}{N}\sum_{i=1}^{N}\ell({\boldsymbol{p}}^{o}_{i},{\boldsymbol{y}}_{i})=-\frac{1}{N}\sum_{i=1}^{N}{\boldsymbol{y}}_{i}^{\top}\log({\boldsymbol{p}}^{o}_{i}),$ (1)
where $\Phi^{o}\triangleq\{\theta^{o},\Theta^{o}\}$ denotes the parameters to be learned, ${\boldsymbol{X}}^{o}\triangleq\{{\boldsymbol{x}}^{o}_{i}\}_{i=1}^{N}$ and $\ell(\cdot)$ denotes the cross entropy loss.
@@ 3.2 Sustained Boosting
By training the model for each modality based on objective function (1), we can obtain multiple individual classifiers. Due to the existence of strong and weak modalities ARM:conf/ijcai/YangYZJ15 , these classifiers exhibit different classification abilities. Hence, we can employ boosting technique GradientBoosting:journal/AS/FriedmanJ to improve the classification ability of weak modality.
@@
Concretely, assuming the classification performance of the $o$-modality requires improvement, we first apply the gradient boosting algorithm to train $n$ classifiers for the $o$-modality. Since feature extraction focuses on common patterns, we set the encoders of all classifiers to be shared. Then the $j$-th classifier can be defined as: $\Phi^{o}_{t}\triangleq\{\theta^{o},\Theta^{o}_{t}\},t\in\{1,\cdots,n\}$. In practice, we adopt multiple fully-connected layers and nonlinear activation rectified linear unit (ReLU) ReLU:journals/corr/abs-1803-08375 to construct our classification module. This module called the configurable classifier, is relatively independent and can be adjusted based on the classification ability. Furthermore, we adopt the shared head structure commonly used in MML MLA:conf/cvpr/ZhangYBY24 ; MVIEW:conf/iccv/0065MW023 ; IGM:conf/ijcai/JiangCY25 to strengthen the interaction between weak and strong modalities during training.
@@
Inspired by gradient boosting GradientBoosting:journal/AS/FriedmanJ , the classification ability can be facilitated through minimizing the residual error introduced by previous classifiers. Concretely, when we learn $t$-th classifier, the residual labels are defined as:
$\hat{\boldsymbol{y}}^{o}_{it}={\boldsymbol{y}}_{i}-\lambda\sum_{j=1}^{t-1}{\boldsymbol{y}}_{i}\odot{\boldsymbol{p}}^{o}_{ij},$
@@
where $\lambda\in[0,1]$ is used to soften hard labels LabelSmoothing:conf/cvpr/SzegedyVISW16 , $\odot$ denotes the element-wise product, and we utilize ${\boldsymbol{y}}_{i}$ to mask non ground-truth labels to ensure the non-negativity of residual labels. Then the objective function can be defined as follows:
$\epsilon({\boldsymbol{x}}^{o}_{i},{\boldsymbol{y}}_{i},t)=\ell\big({\boldsymbol{p}}^{o}_{it},\hat{\boldsymbol{y}}^{o}_{it}\big),$ (2)
@@
where ${\boldsymbol{p}}^{o}_{it}$ denotes the prediction obtained by $t$-th classifier for $i$-th data point. Since we utilize a shared encoder, the encoder will be updated when training the $t$-th classifier. Therefore, other classifiers must be updated simultaneously to prevent performance degradation. The corresponding objective can be formed as:
$\epsilon_{\mathrm{all}}({\boldsymbol{x}}^{o}_{i},{\boldsymbol{y}}_{i},t)$ $=\ell\left({\boldsymbol{p}}^{o}_{it}+\sum_{j=1}^{t-1}{\boldsymbol{p}}^{o}_{ij},{\boldsymbol{y}}_{i}\right)=\ell\left(\sum_{j=1}^{t}{\boldsymbol{p}}^{o}_{ij},{\boldsymbol{y}}_{i}\right).$ (3)
@@
Meanwhile, we have to ensure the first t1t-1 classifiers are well-trained. Hence, we define the following objective for t1t-1 classifiers:
$\epsilon_{\mathrm{pre}}({\boldsymbol{x}}^{o}_{i},{\boldsymbol{y}}_{i},t)$ $=\ell\left(\sum\nolimits_{j=1}^{t-1}{\boldsymbol{p}}^{o}_{ij},{\boldsymbol{y}}_{i}\right).$ (4)
@@
$L({\boldsymbol{x}}^{o}_{i},{\boldsymbol{y}}_{i},t)=\epsilon({\boldsymbol{x}}^{o}_{i},{\boldsymbol{y}}_{i},t)+\epsilon_{\mathrm{all}}({\boldsymbol{x}}^{o}_{i},{\boldsymbol{y}}_{i},t)+\epsilon_{\mathrm{pre}}({\boldsymbol{x}}^{o}_{i},{\boldsymbol{y}}_{i},t).$ (5)
Unlike traditional gradient boosting GradientBoosting:journal/AS/FriedmanJ , our method sustainedly minimizes classification and residual errors by optimizing (5). The loss function of sustained boosting can be formed as:
${\mathcal{L}}_{\mathrm{SUB}}({\boldsymbol{X}}^{o},{\boldsymbol{Y}},n^{o};\Phi^{o})=$ $\frac{1}{N}\sum_{i=1}^{N}L({\boldsymbol{x}}^{o}_{i},{\boldsymbol{y}}_{i},n^{o}),$ (6)
where $n^{o}$ denotes the number of classifier for $o$-modality.
3.3 Adaptive Classifier Assignment
Thus far, we have defined a configurable classifier module and designed a sustained boosting in multimodal learning to enhance the classification performance of weak modality. However, recent studies OGM:conf/cvpr/PengWD0H22 have shown that differences between modalities evolve dynamically due to imbalance issues in multimodal learning. This implies the need to design a strategy for enhancing classification ability that adapts to dynamic changes. Hence, we propose an adaptive classifier assignment strategy to adjust the number of the weak classifier. For simplicity, we redefine the modality classifiers as: $\Phi^{o}_{t}\triangleq\{\theta^{o},\Theta^{o}_{t}\},t\in\{1,\cdots,n^{o}\}$, where $n^{o}$ is the parameter to be updated.
Then, we utilize confident score to monitor the learning status. At $t$-th iteration, confident score can be calculated by:
$\forall o\in\{a,v\},s^{o}_{t}=\frac{1}{N}\sum\nolimits_{i=1}^{N}{\boldsymbol{y}}_{i}^{\top}\left[\sum\nolimits_{j=1}^{n^{o}}{\boldsymbol{p}}^{o}_{ij}\right].$
The confident score reflects the classification ability of the models. Hence, if $s^{a}_{t}-\sigma s^{v}_{t}>\tau$, we assign a new configurable classifier for video modality at this iteration, where $\sigma\geq 1$ is the coefficient. $\tau$ is the dead zone for fault tolerance. On the contrary, we also assign a new configurable classifier for audio modality if $s^{a}_{t}-\sigma s^{v}_{t}<\tau$.
Our algorithm is summarized in Algorithm (1). In practice, we perform adaptive classification assignment strategy to determine if we need to adjust the classification ability every $t_{N}$ iterations.
Algorithm 1 Learning algorithm of our proposed method.
1:Training data $\boldsymbol{X}$}, category labels $\boldsymbol{Y}$}.
2:The learned DNN models for all modalities.
3:INIT Initialize the number of classifier $n^{a}=1$, $n^{v}=1$. Initialize iteration $t=1$. Initialize DNN parameters $\Phi^{a}_{t}$ and $\Phi^{v}_{t}$.
4:for $t=1\mapsto\#iterations$ do
5:  Sample a mini-batch ${\boldsymbol{X}}_{t}=\{({\boldsymbol{x}}_{i}^{a},{\boldsymbol{x}}_{i}^{v})\}_{i=1}^{n_{b}}$; $\triangleright$ Learn MML models.
6:  $\forall{\boldsymbol{x}}_{i}^{a},{\boldsymbol{x}}_{i}^{v}\in{\boldsymbol{X}}_{t}$, calculate features ${\boldsymbol{u}}_{i}^{a}$ and ${\boldsymbol{u}}_{i}^{v}$;
7:  Calculate predictions $\{{\boldsymbol{p}}^{a}_{ij}\}_{j=1}^{n^{a}}$ and $\{{\boldsymbol{p}}^{v}_{ij}\}_{j=1}^{n^{v}}$.
8:  Calculate loss in (6) based on predictions;
9:  Update DNN parameters $\Phi^{a}_{t}$ and $\Phi^{v}_{t}$ based on SGD;